Need the R Studio Script for the following excercises (R Studio only)



P1. According to the 2001 Youth Risk Behavior Surveillance by the Centers for Disease Control and Prevention, 39% of the 10th-graders surveyed said that they watch three or more hours of television on a typical school day. Assume that this percentage is true for the current population of all 10th-graders. What is the probability that 86 or more of the 10th-graders in a random sample of 200 watch three or more hours of television on a typical school day? The required probability .5 . 3621 .1379 78 85.5 0 1.09 P2. After a general culture test, it is observed that the scores obtained follow a distribution with an N distribution (65, 18). It is desired to classify the examinees into three groups (low general culture, acceptable general culture, excellent general culture) so that there is in the first 20% population, 65% the second and 15% in the third. what should be the scores that mark the passage from one group to the other? P3. Several intelligence tests gave a score that follows a normal law with mean 100 and standard deviation 15. a) Determine the percentage of the population that would obtain a coefficient between 95 and 110. b) In a population of 2500 individuals, how many individuals are expected to have a coefficient greater than 125
media%2F677%2F6778642b-3a71-42a9-bf62-52
media%2F1cf%2F1cf28359-dabb-4f04-a54d-ea
Answer #1

Solution:-

P1) The probability that 86 or more of the 10th graders in a random sample of 200 watch three or more hours of television on a typical school day is 0.123.

p = 0.39

n = 200

z = \frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}}

z = 1.16

P(z > 1.16) = 0.123